### v2

# Advancing GPU IPC for Stiff Affine-Deformable Simulation

Abstract.
Incremental Potential Contact (IPC) is a widely used, robust, and accurate method for simulating complex frictional contact behaviors. However, achieving high efficiency remains a major challenge, particularly as material stiffness increases, which leads to slower Preconditioned Conjugate Gradient (PCG) convergence, even with the state-of-the-art preconditioners. In this paper, we propose a fully GPU-optimized IPC simulation framework capable of handling materials across a wide range of stiffnesses, delivering consistent high performance and scalability with up to 10$\times$ speedup over state-of-the-art GPU IPC methods. Our framework introduces three key innovations: 1) A novel connectivity-enhanced Multilevel Additive Schwarz (MAS) preconditioner on the GPU, designed to efficiently capture both stiff and soft elastodynamics and improve PCG convergence at a reduced preconditioning cost. 2) A $C^{2}$-continuous cubic energy with an analytic eigensystem for strain limiting, enabling more parallel-friendly simulations of stiff membranes, such as cloth, without membrane locking. 3) For extremely stiff behaviors where elastic waves are barely visible, we employ affine body dynamics (ABD) with a hash-based multi-layer reduction strategy for fast Hessian assembly and efficient affine-deformable coupling. We conduct extensive performance analyses and benchmark studies to compare our framework against state-of-the-art methods and alternative design choices. Our system consistently delivers the fastest performance across soft, stiff, and hybrid simulation scenarios, even in cases with high resolution, large deformations, and high-speed impacts. Our framework will be fully open-sourced upon acceptance.
GPU Programming, Incremental Potential Contact, Elastodynamics, Finite Element Method, Affine Body Dynamics, Preconditioning, Cloth Simulation
ccs: Computing methodologies Physical simulationccs: Computing methodologies Parallel algorithms
Figure 1. London bus. We model a $0.7m\times 1.2m\times 2.3m$ bus, including its frame, gears, and wheels, based on the puzzle instructions from Wooden City®. Despite the eye-norm modeling accuracy, our simulation robustly captures the realistic non-smooth motion caused by potential interference in the low-precision gear system and the self-adjustment of motor speed (averaging $3m/s$) relative to resistance (see our supplemental video). The simulation includes 265K surface triangles and an average of 118K rapidly changing contact pairs, achieving efficient performance at $1.56s$ per time step ($\Delta t=5ms$).
1. Introduction
Incremental Potential Contact (IPC) (Li et al., 2020a) is a cutting-edge elastodynamic contact simulation method widely used in computer graphics, computational mechanics, robotics, etc. Despite its robustness, accuracy, and differentiability in simulating complex frictional contact behaviors, IPC’s efficiency remains a significant bottleneck, limiting its full potential. Several variants have been proposed to address IPC’s efficiency issues, often at the expense of accuracy (Lan et al., 2021, 2023; Li et al., 2023; Lan et al., 2022b).
To accelerate IPC without sacrificing accuracy, Huang et al. (2024) introduced GIPC with a GPU-friendly redesign of the numerical algorithms. This included replacing direct factorization with a Preconditioned Conjugate Gradient (PCG) solver and proposing a Gauss-Newton approximation of the barrier Hessian matrices with analytic eigensystems. While GIPC is effective, its efficiency deteriorates significantly as object stiffness increases. This is mainly due to the growing condition number of the global linear system, which requires more PCG iterations to solve. To improve the PCG convergence of stiff material, Wu et al. (2022) proposed the multilevel additive Schwarz (MAS) preconditioner (Wu et al., 2022). Their approach involves sorting the nodes based on Morton codes and building a hierarchy by grouping the nodes at each level. Despite its effectiveness, the lack of consideration for mesh connectivity during reordering leads to suboptimal domain hierarchy construction. This results in high construction costs, additional overhead for deformable simulations and challenges for GPU optimization.
When simulating stiff elastic thin shells like cloth, another challenge arises, which is membrane locking. With linear triangle elements, the stiff membrane energy (Young’s modulus around $10MPa$ for cloth (Penava et al., 2014)) will often result in nonnegligible extra artificial bending resistance, forming sharp creases and plastic appearances in the simulation results (Figure 12 top). Simulating cloth with smaller stiffness can result in more realistic wrinkles, but it will suffer from over-elongation issues. To tackle this challenge, Li et al. (2021) propose to augment soft membrane energy with a barrier-based strain-limiting term to prevent cloth from over-stretching while avoiding membrane locking. This strategy enables realistic cloth simulation within the IPC framework, but the required exact strain limit satisfaction necessitates a backtracking-based line search filtering scheme, as the updated strain has a complicated relation to the step size, making analytic expressions unavailable. Additionally, numerical eigendecomposition is needed for computing a positive semi-definite approximation of the strain-limiting term’s Hessian matrix, which further complicates GPU optimization.
For even stiffer problems where elastic waves are barely visible, objects can be treated as rigid (Ferguson et al., 2021) or stiff affine (Lan et al., 2022a) bodies in the IPC framework for a reduced number of degrees of freedom (DOF). But to accurately simulate contact behaviors, surface elements from the original input geometry are used, which also makes simulating rigid-deformable coupled scenarios convenient. Chen et al. (2022) introduced a unified Newton barrier method for stiff affine-deformable simulation, possibly with articulation constraints. However, although some components are GPU-accelerated, the primary simulation processes still execute on the CPU, leading to suboptimal performance. ZeMa (Du et al., 2024) is another GPU IPC framework for stiff affine-deformable simulation, with most processes parallelized on the GPU, except for the linear system, which is solved on the CPU using a direct solver. However, ZeMa lacks a well-optimized contact Hessian assembly algorithm, as it accumulates the $12\times 12$ dense contact Hessian matrices to the affine body DOFs atomically, where conflicting operations can significantly impede the performance, especially when there are a large number of contacts. Moreover, direct solvers often fall short in large-scale simulations.
In summary, there are still plenty of rooms for optimizing linear solver preconditioners, strain limiting, global Hessian matrix assembly, etc., for realizing a highly GPU-optimized IPC framework that can efficiently simulate large-scale affine-deformable coupled scenarios. In this paper, we propose such a framework, achieving up to 10×10$\times$10 × speedup compared to GIPC via the following 3 major innovations:
  • A novel connectivity-enhanced MAS preconditioner on the GPU that achieves improved PCG convergence at a lower precomputation and per-iteration cost (section 4). Our preconditioner consistently performs effective and well-structured aggregations, which supports smaller blocksizes and further GPU optimizations based on warp reduction.
  • A $C^{2}$-continuous cubic strain-limiting energy with an analytic eigensystem, enabling realistic cloth simulation without membrane locking (section 5). As numerical eigendecomposition and line search filtering for the feasibility of the strain limits are not needed, our model supports highly GPU-parallelized computations.
  • A hash-based multi-layer reduction strategy for fast Hessian matrix assembly (section 6). Our strategy significantly reduces the number of numerical operations, and it enables the development of a memory-efficient symmetric blockwise sparse matrix-vector multiplication method to further boost PCG performance.
In section 7, we perform extensive and rigorous performance analyses and benchmark studies to validate our framework and compare it to state-of-the-art GPU IPC systems and alternative design choices that may seem reasonable but suffer from suboptimal performance in practice. Our framework exhibits the fastest performance in soft, stiff, and hybrid simulation scenarios, even with high resolution, extreme deformation, and high-speed impacts. Our system will be fully open-sourced upon acceptance.
2. Related Works
Contact Simulation
Simulating frictional contact for (nearly) rigid and deformable solids has been an extensively studied topic in both computer graphics and computational mechanics. Starting from a few decades ago, various methods have been developed, ranging from impulse-based methods (Bridson et al., 2002; Harmon et al., 2009), impact zone methods (Harmon et al., 2008; Tang et al., 2018; Li et al., 2020b), and more constraint-based methods (Kaufman et al., 2008; Otaduy et al., 2009; Macklin et al., 2019; Verschoor and Jalba, 2019; Allard et al., 2010) to fictitious domain methods (Jiang et al., 2017; Müller et al., 2015; Misztal and Bærentzen, 2012; Pagano and Alart, 2008), etc. These methods are effective and efficient in many situations, but they lack guarantees on algorithmic convergence and generating penetration-free results, especially when simulating challenging examples where extensive parameter tuning is often also required when varying set-ups. We refer to Andrews et al. (2022) for a comprehensive review.
More recently, Li et al. (2020a) proposed Incremental Potential Contact (IPC), which simulates the nonsmooth frictional contact behaviors of deformable solids by approximating them using smooth constitutive models with bounded error. Equipped with a filter line search scheme and the projected Newton method, penetration-free results and algorithmic convergence are guaranteed within an optimization time integrator. IPC has shown effectiveness in various application scenarios, including cloth reconstruction (Zheng et al., 2024), material modeling of interlocked rigid components (Tang et al., 2023), multi-material coupled simulations (Xie et al., 2023; Jiang et al., 2022; Li et al., 2024), etc. Despite being robust and accurate, improving the efficiency of IPC while maintaining its reliability is still a major challenge.
Accelerating IPC Performance
Over the past few years, several methods have been proposed to enhance IPC’s efficiency. Lan et al. (2021) reduced the DOFs of deformable bodies to their medial axis and derived contact and friction forces on the associated slab primitives. With significantly fewer DOFs, the efficiency improved substantially at the expense of accuracy, and penetration-free results were only guaranteed on their medial representation. Lan et al. (2022b) incorporated IPC into the projective dynamics framework on the GPU, delaying contact constraint set updates and continuous collision detection (CCD) to once per set of inner iterations to efficiently generate penetration-free results. However, their spring-based approximation to the barrier-based contact model limited their performance in challenging scenarios. Lan et al. (2023) applied a stencil-wise coordinate descent method to solve the IPC system. With a penetration-free warm start and graph coloring-based GPU parallelization, their method achieved fast performance. However, as a block coordinate descent method, its performance drops quickly with stiff materials or a large number of contacts. Wang et al. (2023a) and Li et al. (2023) both applied time-splitting techniques to decouple contact and elasticity simulation, achieving faster performance but requiring small time steps for stiff problems due to the lack of unconditional stability. These methods trade accuracy for efficiency using simplified models or specialized strategies, often introducing extra parameters for tuning.
More recently, Huang et al. (2024) introduced GIPC, a fully GPU-optimized IPC method with an inexact Gauss-Newton solver using the MAS preconditioner (Wu et al., 2022). GIPC also derived approximated contact energy Hessians with analytic eigensystems, avoiding GPU-unfriendly numerical eigendecomposition. Concurrently, Du et al. (2024) proposed ZeMa, another GPU IPC framework supporting the coupling of deformable bodies with stiff affine bodies. Both GIPC and ZeMa achieved significant speed-ups without sacrificing accuracy. However, as discussed earlier, their high performance could unavoidably degrade when there are stiff materials or a large number of contacts. We thus follow this path to keep advancing GPU IPC for stiff elastodynamics without sacrificing accuracy and robustness.
Multiresolution Methods
One of the most critical design choices on the GPU is to apply iterative solvers instead of direct solvers for linear systems. Multigrid is a popular GPU-friendly preconditioner often used in iterative linear solvers, such as the PCG method, for efficiently solving ill-conditioned systems. The major challenge in multigrid methods is the development of effective restriction and prolongation operators, especially for unstructured meshes. Researchers in computer graphics have explored various geometric multigrid methods (GMG) (Wang et al., 2020, 2018; Xian et al., 2019; Zhu et al., 2010) by constructing mesh hierarchies, where numerous implementation issues are often encountered. Compared to GMG, another type of multigird, the algebraic multigrid (AMG) methods (Naumov et al., 2015; Demidov, 2019; Tamstorf et al., 2015), do not require explicitly constructing coarser geometries. However, they often require a significant amount of precomputation to analyze the structure of the matrix, making them more expensive when handling problems with varying matrix structures. This inspires Wu et al. (2022) to propose a multilevel additive schwarz (MAS) (Dryja and Widlund, 1990) preconditioner, incorporating the idea of domain decomposition to achieve fast convergence with low per-iteration costs when applied to PCG. But its node reordering based on Morton code is not aware of the mesh connectivity, which may lead to suboptimal aggregation. We thus propose a connectivity-enhanced MAS based on METIS (Karypis and Kumar, 1998) to tackle this issue and achieve even faster performance.
Strain Limiting and Eigenanalysis
To realistically simulate cloth using linear elements, strain limiting is often applied to avoid membrane locking. Most existing methods (Narain et al., 2012; Wang et al., 2010; Thomaszewski et al., 2009; English and Bridson, 2008; Goldenthal et al., 2007) enforce a bound constraint on the strain measure per element using augmented Lagrangian approaches or post-projection techniques. While these methods do not guarantee strict satisfaction of strain limits, they can effectively simulate membrane locking-free behaviors. However, when coupled with frictional contact, these methods often introduce artifacts due to the need for smaller time steps or difficulties in achieving solver convergence. More recently, Li et al. (2021) applied a barrier method to handle strain-limiting constraints, guaranteeing exact constraint satisfaction in a monolithic manner. However, it requires expensive backtracking-based line search filtering and numerical eigendecomposition, which are both GPU-unfriendly. Here, the backtracking line search is responsible for finding configurations that are both feasible and with a smaller objective function. With the development of analytical eigenanalysis on distortion energies, such as the isotropic (Smith et al., 2019; Lin et al., 2022) and anisotropic (Kim et al., 2019) elasticity, membrane (Panetta, 2020; Kim, 2020), bending (Wang et al., 2023b; Wu and Kim, 2023), and contact (Shi and Kim, 2023; Huang et al., 2024) energies, they have become popular components of modern GPU simulation methods. We thus propose a cubic strain-limiting energy with analytic eigensystems so that realistic cloth dynamics can be more efficiently simulated without the need of numerical eigendecomposition and backtracking-based line search filtering.
3. Background and preliminaries
Figure 2. An overview of 1 Newton iteration in our unified GPU IPC framework. Our simulation focuses on cloth, soft bodies with varying stiffness, and stiff affine bodies (Lan et al., 2022a). Cloth and soft bodies are modeled using FEM. Contact between all objects is handled using IPC (Li et al., 2020a). We solve the system with a Newton-PCG solver and use Continuous Collision Detection (CCD) to filter the search direction and prevent penetration, applying backtracking line search to ensure energy decrease. Our contributions, highlighted in light pink blocks, include a novel strain limiting energy with an analytic eigensystem for efficient PSD projection, an optimized linear solver with a connectivity-enhanced MAS preconditioner, and a highly-parallelized global Hessian assembly strategy for affine-deformable coupling.
3.1. Unified Incremental Potential
IPC (Li et al., 2020a) formulates implicit time integration of elastodynamic contact as a minimization problem:
(1) $\mathbf{x}^{t+\Delta{t}}\approxeq\operatorname*{arg\,min}_{\mathbf{x}\in\rm I\!R^{3n}}E_{s}(\mathbf{x}),$
where $\mathbf{x}$ is the world-space positions of the $n$ nodes, followed by a velocity update $\mathbf{v}^{t+\Delta{t}}=(\mathbf{x}^{t+\Delta{t}}-\mathbf{x}^{t})/\Delta t$, taking implicit Euler as an example. The objective function
(2) $E_{s}(\mathbf{x})=\frac{1}{2}(\mathbf{x}-\mathbf{\hat{\mathbf{x}}})^{T}\mathbf{M}_{s}(\mathbf{x}-\mathbf{\hat{\mathbf{x}}})+\Delta t^{2}\Psi_{s}(\mathbf{x})+B(\mathbf{x})+D(\mathbf{x})$
is the Incremental Potential (IP), where $\hat{\mathbf{x}}=\mathbf{x}^{t}+\Delta t\mathbf{v}^{t}+\Delta t^{2}\mathbf{M}_{s}^{-1}\mathbf{f}_{e}$, $\mathbf{M}_{s}$ is the mass matrix, $\mathbf{f}_{e}$ is the external force, $B$ is the contact barrier potential, and $D$ is the approximated friction potential. The total elasticity energy $\Psi_{s}(\mathbf{x})=\Psi_{vol}(\mathbf{x})+\Psi_{memb}(\mathbf{x})+\Psi_{bend}(\mathbf{x})+\Psi_{strain}(\mathbf{x})$ contains volumetric strain energies ($\Psi_{vol}(\mathbf{x})$) as well as membrane ($\Psi_{memb}(\mathbf{x})$), bending ($\Psi_{bend}(\mathbf{x})$), and strain-limiting ($\Psi_{strain}(\mathbf{x})$) energies for thin shells.
The IP for simulating affine bodies (Lan et al., 2022a) is defined as
(3) $E_{r}(\mathbf{q})=\frac{1}{2}(\mathbf{q}-\mathbf{\hat{\mathbf{q}}})^{T}\mathbf{M}_{r}(\mathbf{q}-\mathbf{\hat{\mathbf{q}}})+\Delta t^{2}\Psi_{r}(\mathbf{q})+B(\mathbf{x}(\mathbf{q}))+D(\mathbf{x}(\mathbf{q})),$
where $\mathbf{q}\in\rm I\!R^{12\alpha}$ is the reduced space coordinates of the $\alpha$ affine bodies. For affine body $j$, $\mathbf{q}_{j}=[\mathbf{p}_{j}^{T},\mathbf{A}_{j1}^{T},\mathbf{A}_{j2}^{T},\mathbf{A}_{j3}^{T}]^{T}$, with $\mathbf{p}_{j}\in\rm I\!R^{3}$ the translation vector and $\mathbf{A}_{j}\in\rm I\!R^{3\times 3}$ the affine deformation matrix. Here, $\mathbf{x}_{i}(\mathbf{q}_{j})=\mathbf{A}_{j}\bar{\mathbf{x}}_{i}+\mathbf{p}_{j}$ is the current position of full-space node $i$ on affine body $j$ for measuring contact energies, with $\bar{\mathbf{x}}_{i}$ the rest-state full space coordinates. The Jacobian matrix $\mathbf{J}_{ij}=\frac{\partial\mathbf{x}_{i}}{\partial\mathbf{q}_{j}}\in\rm I\!R^{3\times 12}$ maps information between the reduced and full spaces. For example, the mass matrix $\mathbf{M}_{r}$ of affine body $j$ is calculated as $\mathbf{M}_{r}=\sum_{i}m_{i}\mathbf{J}_{ij}^{T}\mathbf{J}_{ij},$ where $m_{i}$ is mass of each node $i$. Similarly, the contact gradient of affine body $j$ is calculated as $\nabla_{\mathbf{q}_{j}}B=\sum_{i}\mathbf{J}_{ij}^{T}\nabla_{\mathbf{x}_{i}}B.$ From this, we can see that computing contact forces and Hessian matrices for the reduced DOFs could require accumulating values from a significant amount of contact pairs, which can be expensive.
Combining Equation 2 and Equation 3, denoting nodal positions of unreduced deformable solids as $\mathbf{x}_{s}$, the unified affine-deformable coupled IP (Chen et al., 2022) is defined as
$E(\mathbf{q},\mathbf{x}_{s})$ $=\frac{1}{2}(\mathbf{x}_{s}-\mathbf{\hat{\mathbf{x}}}_{s})^{T}\mathbf{M}_{s}(\mathbf{x}_{s}-\mathbf{\hat{\mathbf{x}}}_{s})+\Delta t^{2}\Psi_{s}(\mathbf{x}_{s})$
$\,\,+\frac{1}{2}(\mathbf{q}-\mathbf{\hat{\mathbf{q}}})^{T}\mathbf{M}_{r}(\mathbf{q}-\mathbf{\hat{\mathbf{q}}})+\Delta t^{2}\Psi_{r}(\mathbf{q})$
(4) $\,\,+B([\mathbf{x}_{s}^{T},\mathbf{x}(\mathbf{q})^{T}]^{T})+D([\mathbf{x}_{s}^{T},\mathbf{x}(\mathbf{q})^{T}]^{T}),$
and we minimize this energy w.r.t. $\{\mathbf{q},\mathbf{x}_{s}\}$ per time step to simulate the coupled system:
(5) $(\mathbf{q},\mathbf{x}_{s})^{t+\Delta{t}}\approxeq\operatorname*{arg\,min}_{\mathbf{x}_{s}\in\rm I\!R^{3n},\mathbf{q}\in\rm I\!R^{12\alpha}}E(\mathbf{q},\mathbf{x}_{s}).$
To accelerate local Hessian computation on the GPU, we utilize potential energies with analytic eigensystems (subsection 3.2). Specifically, for collision energy $B$ and friction energy $D$, we employ the eigenanalysis from GIPC (Huang et al., 2024). The Stable Neo-Hookean model (Smith et al., 2018) is applied for $\Psi_{vol}$, while the bending energy from (Wu and Kim, 2023) is used for $\Psi_{bend}$. To efficiently assemble $\Psi_{memb}$ and $\Psi_{strain}$, we use Kim (2020)’s membrane energy and propose a cubic strain-limiting energy with analytic eigensystem, detailed in section 5.
3.2. IP Minimization and Linear Solves
To minimize $E(\mathbf{u})$, where $\mathbf{u}=\{\mathbf{q},\mathbf{x}\}$, line search methods are often used, which iteratively search along a descent direction by minimizing a local quadratic proxy of the IP in iteration $i$:
(6) $E_{i}(\mathbf{u})=E(\mathbf{u}_{i})+\left(\mathbf{u}-\mathbf{u}_{i}\right)^{T}\nabla E(\mathbf{u}_{i})+\frac{1}{2}\left(\mathbf{u}-\mathbf{u}_{i}\right)^{T}\mathbf{P}(\mathbf{u}_{i})\left(\mathbf{u}-\mathbf{u}_{i}\right).$
The descent direction is obtained by solving the linear system $\mathbf{P}(\mathbf{u}_{i})\mathbf{d}=-\nabla E(\mathbf{u}_{i})$, where $\mathbf{d}=\mathbf{u}_{i}^{*}-\mathbf{u}_{i}$ and $\mathbf{u}_{i}^{*}=\operatorname*{arg\,min}_{\mathbf{u}}E_{i}(\mathbf{u})$. Here, $\mathbf{P}$ is a symmetric positive definite (SPD) proxy matrix approximating $\nabla^{2}E$ for fast convergence, with projected Newton (Li et al., 2020a) or Gauss-Newton (Huang et al., 2024) approximations shown effective. The new iterate is computed as $\mathbf{u}_{i+1}=\mathbf{u}_{i}+\alpha\mathbf{d}$, where $\alpha$ is the step size calculated via backtracking line search, with filtering to ensure feasibility imposed by contact and strain limits (Li et al., 2021). The minimization terminates when the iterate is sufficiently close to the local minimum, e.g. when $\|\mathbf{d}\|\leq\varepsilon_{d}$, where $\varepsilon_{d}$ is the Newton tolerance.
To solve the linear system $\mathbf{P}(\mathbf{u}_{i})\mathbf{d}=-\nabla E(\mathbf{u}_{i})$, Cholesky factorization, using e.g. CHOLMOD (Chen et al., 2008), is a popular solution. These direct solvers perform well on ill-conditioned systems but their performance decreases with increasing DOFs, especially for large-scale simulations. Additionally, the high memory cost of direct solvers results in worse performance on the GPU compared to the CPU (Lan et al., 2021). Conversely, iterative solvers, such as the preconditioned conjugate gradient (PCG) method, iteratively approach the solution via matrix-vector multiplications (SpMV), making them more practical for large-scale simulations.
Although PCG has superior convergence rate compared to many non-Krylov iterative solvers, such as Jacobi and Gauss-Seidel, it still requires many iterations to reach low error when the linear system is ill-conditioned. An effective preconditioner that approximates the inverse of the proxy matrix can make the system better conditioned and easier to solve. MAS (Wu et al., 2022) is currently the best option considering both convergence speed and the overhead of computing and applying the preconditioner. Additionally, the high computational cost of matrix assembly and SpMV in each iteration often makes PCG the major bottleneck in simulation. Thus, designing GPU-parallel algorithms for PCG is also critical. See Figure 2 for an overview of our simulation framework.
Figure 3. A node sorting example using Morton code. Red disks represent nodes, green lines show topology connections, and dotted blue lines indicate Morton code order. After sorting, nodes will follow the character sequence shown. Note that this is a contrived example for clarity.
4. Connectivity-Enhanced MAS Preconditioner with GPU Optimization
The multilevel additive Schwarz (MAS) preconditioner (Wu et al., 2022) enables efficient linear solver convergence with excellent scalability w.r.t. problem size and stiffness in elastodynamics problems. However, its node reordering has some limitations, which makes it less effective in cases with complex geometry or softer materials. We start by a comprehensive analysis of these limitations (subsection 4.1), and then propose a connectivity-enhanced MAS construction (subsection 4.2) together with further GPU optimizations (subsection 4.3) to achieve a $\times2$ faster performance in average compared to Wu et al. (2022).
4.1. Discussions of MAS
MAS constructs the domain hierarchy by first sorting all mesh vertices using Morton codes, and then grouping these vertices into subdomains of size $N$. Next, connected vertices within each subdomain are merged to create the next-level simulation domain. These merged vertices, called super nodes, are then grouped again to form the domains at subsequent levels until a maximum level is reached, no further merges could be performed, or only a single subdomain is present. During preconditioning, the inverse of each subdomain’s Hessian matrix is multiplied with the mapped input vector, and the resulting vectors are mapped and summed to produce the output. With a fixed number of levels, maximizing node connectivity within each subdomain is critical to making dimensionality reduction more effective, which allows information to propagate farther at coarser levels and ensures faster convergence. See our supplemental document for more details.
We identify two major areas for potential improvement in MAS: 1) The convergence of MAS heavily relies on node reordering using Morton codes. While Morton codes capture spatial information, they do not consider mesh connectivity, which can result in slower convergence. 2) To obtain denser node connectivity within each subdomain and improve aggregation efficiency, a subdomain size of 32 vertices is typically used in MAS. However, this leads to significant overhead in terms of precomputation and preconditioning, especially when using double-precision floating-point numbers. Reducing the subdomain size to 16 vertices can decrease computational costs, but it may also slow down convergence, as fewer nodes are merged within each subdomain, potentially resulting in worse overall performance (Wu et al., 2022).
Figure 4. MAS aggregation of Figure 3 using Morton code. Within each subdomain (size 4 here, separated by dashed blue lines), connected nodes are merged into a super node (orange block) at the next coarser level. The order of the super node is determined by the lowest order of the merged nodes. For example, in level 1, nodes $b$ and $d$ are merged and positioned before $c$ because $b$ has a lower order. The subdomains are then reconstructed from sets of 4 consecutive super nodes, and here this process continues until no further merging is possible at a certain level.
Case Study
Let’s examine the first issue more closely with an example (Figure 3). In this case, we assume a domain size of 4 for MAS, meaning the nodes are grouped into subdomains $\Omega=\Omega_{0}\cup\Omega_{1}\cup\Omega_{2}\cup\Omega_{3}$, each with 4 nodes, at level 0, as shown in Figure 4. In $\Omega_{0}=\{a,b,c,d\}$, only nodes $\{b,d\}$ are connected. Consequently, at the coarser level 1, these two nodes are merged into a super node. Similarly, nodes $\{f,h\}$, $\{i,j,k,l\}$, $\{m,o\}$, and $\{n,p\}$ are merged at level 1. However, starting from level 1, there is no connectivity between any super nodes in the first subdomain, and only a few can be merged in the other subdomains. This results in 3 levels for these 16 nodes, with inefficient aggregation, leading to poor performance of the MAS preconditioner. In fact, the performance can even be worse than that of a simple block Jacobi preconditioner in practice.
Figure 5. Our connectivity-enhanced MAS aggregation of Figure 3. The orange blocks represent super nodes at the coarser levels. At level 0, our method groups nodes based on mesh connectivity, and so each group becomes 1 super node at level 1, completing the aggregation.
Since the aggregation is based on node connectivity, a natural improvement is to group the nodes directly by their connectivity at level 0, ensuring that the 4 nodes in each subdomain are fully connected and can be merged into a single super node, as shown in Figure 5. In this way, at level 1, there will be only 4 super nodes that cover all the original nodes in the mesh, allowing us to terminate the aggregation with just one subdomain. As a result, this method constructs an efficient two-level MAS that can be used to effectively approximate the global Hessian.
4.2. Connectivity-Enhanced MAS Construction
The example above demonstrates the importance of considering mesh connectivity during node reordering for optimal node aggregation. However, in practice, achieving perfect aggregation is both expensive and challenging. Therefore, we propose a practical connectivity-enhanced strategy that efficiently approximates optimal aggregation. Our core idea is to use METIS (Karypis and Kumar, 1998) to partition the mesh nodes before the simulation and then reorder them with padding at level 0, optimizing node connectivity within each subdomain.
METIS is a fast graph partitioning method that can divide mesh nodes into a specific number of partitions, minimizing inter-partition connectivity and often leading to dense connectivity within each partition. However, METIS does not guarantee that each partition will have exactly the same number of nodes, as achieving perfectly balanced partitions is expensive and sometimes impossible. Handling blocks of varying sizes on the GPU can degrade performance and complicate implementation. Thus, we can treat METIS partitioning as a reordering step, and then distribute all nodes evenly into each MAS subdomain. However, this can still result in suboptimal aggregation since the nodes in the same subdomain may come from different METIS partitions as shown in 6(a).
Figure 6. Aligning METIS partitions and MAS subdomains. The red blocks represent subdomains from the METIS partition, the deep blue rectangle shows the MAS subdomain, and the green-blue blocks represent inserted zero entries as padding. (a) An example where the METIS partition sizes are 3 and 4, while the MAS subdomain size is 4, leading to isolated nodes within subdomains and potentially suboptimal aggregation (subsection 4.1). (b) Our solution inserts zero entries to prevent isolated nodes caused by misalignment between the METIS partition and MAS subdomains.
To address this issue and ensure effective aggregation, we propose an index mapping scheme to assign each partition to a separate MAS subdomain and pad the empty entries as illustrated in 6(b). We first set the number of METIS partitions $M$ as
(7) $M$ $=\left\lceil\frac{V}{N-N_{o}}\right\rceil,$
[…]
To implement this approach, we first construct a mapping array of size $N\cdot M$, which maps each node to a slot in its corresponding MAS subdomain, as shown in Figure 7. This mapping array is then used to build the MAS preconditioner. To ensure correct data access during construction, we also create a remapping array, which helps GPU threads access the correct node data for each entry in the mapping array. Padded entries, where no nodes are mapped, are marked with the value $-1$ (see Figure 7). The remaining MAS levels are constructed using the original method based on node connectivity. By updating the mapping per time step considering the connectivity introduced by the current contact pairs, our approach reduces connectivity issues and enables more effective aggregation, improving the convergence speed of the MAS-preconditioned linear solver by up to 2.4$\times$.
Figure 7. Our mapping and remapping scheme. Nodes in each METIS partition are mapped to a separate MAS subdomain using a mapping array (stored in contiguous memory, aligned here for clarity) with padding shown in blue. Our remapping array maps MAS subdomain nodes to the corresponding mesh nodes for easy access to geometric data.
Another significant advantage of our method is the ability to use smaller subdomain sizes, resulting in substantial performance gains when computing the inverse of diagonal blocks and applying the preconditioner. In our implementation, we use a subdomain size of 16, compared to the size 32 used in (Wu et al., 2022). Despite the smaller size, our method remains effective as node connectivity within each subdomain is optimized. This enables cost-effective preconditioning while maintaining a high convergence speed, achieving up to 3$\times$ speedup in the PCG solve.
Algorithm 1 Compute MAS preconditioner
1:Input: Global Hessian matrix $\mathbf{AG}$triangleright an array of tuples: [block row index, block column index, $3\times 3$ block]
2:Output: MAS preconditioner $\mathbf{M}_{MAS}^{-1}$
3:Method:
4:for  $GlobalThreadId$ $=$ 00, $1$, …, $len(\mathbf{AG})-1$ do in parallel
5:     $[r,c,\mathbf{M3}]=\mathbf{AG}(GlobalThreadId)$triangleright get $3\times 3$ block
6:     $[domain\_Idx,domain\_Idy]=\mathbf{Map}(r,c)$triangleright align METIS and MAS subdomain with padding (Figure 6)
7:     $m=domain\_Idy/N$triangleright get MAS subdomian id
8:     $l=0$triangleright MAS level, starting from 0
9:     if $domain\_Idx/N==domain\_Idy/N$ then $\triangleright$ if nodes $r$ and $c$ belong to the same level-0 MAS subdomain, write $\mathbf{M3}$ to the level-0 MAS matrix
10:         $\mathbf{M}_{0,MAS}(m)[domain\_Idx\%N,domain\_Idy\%N]=\mathbf{M3}$
11:     $\triangleright$ add $\mathbf{M3}$ to higher level MAS matrix if needed
12:         for $l=1,2,...,levels-1$ do
13:              if $r$ and $c$ in the same subdomain at level $l$ then
14:                  AtomicAdd $\mathbf{M3}$ to $\mathbf{M}_{l,MAS}$
15:              end if
16:         end for
17:     end if
18:end for
19:for $BId$ $=$ 00, $1$, …, $len(\mathbf{M}_{0,MAS})-1$ do in parallel $\triangleright$ accumulate all $3\times 3$ blocks in $\mathbf{M}_{0,MAS}$ by warp reduction and add it to all higher-level MAS matrices
20:     $\mathbf{M3}={warpReduction}(\mathbf{M}_{0,MAS}(BId))$
21:     for $l=1,2,...,levels-1$ do
22:         AtomicAdd $\mathbf{M3}$ to $\mathbf{M}_{l,MAS}$
23:     end for
24:end for
25:return $\mathbf{M}_{MAS}^{-1}$
4.3. GPU Warp Reduction for Connectivity-Enhanced MAS
Our connectivity-enhanced MAS construction ensures that the nodes within each subdomain at level 0 are densely connected. This property opens up new opportunities for additional GPU optimizations in efficiently computing the proxy Hessian matrix at coarser MAS levels.
In GPU MAS (Wu et al., 2022), since nodes in each subdomain can be mapped to an arbitrary number of super nodes at the next coarser level, it is nontrivial to design an efficient reduction scheme for accumulating the matrix entries to assemble the proxy matrix at the next level. Instead, they construct a mapping array to establish connections between mesh nodes and super nodes at different levels, and then apply atomic operations to compute the proxy matrices at higher levels. Take Figure 4 as an example, node $\{f\}$ in level 0 maps to $\{f,h\}$ in level 1 and 2, and then to $\{f,h,n,p\}$ in level 3. The data associated with $\{f\}$ will be added to the super nodes at coarser levels via atomic operations. However, as we see, as the structure becomes coarser, more atomic additions of a larger number of matrix entries are needed, which will result in more writing conflicts and increase computational cost, especially when using double-precision floating-point numbers.
In our case, the nodes within the same level-0 subdomain map to a small number (often the same) super nodes in coarser levels. This allows for the application of a warp reduction method to accumulate the level-0 matrix blocks. The accumulated matrix is then used to construct the sub-matrix of the next coarser levels using atomic operations. This approach effectively reduces the cost of MAS matrix precomputation, since significantly less atomic operations are needed as the level-1 matrix can be efficiently computed using warp reduction. This also allows for more levels in the MAS preconditioner with lower computational cost. See Algorithm 1 and our supplemental document for more details.
5. Cubic Strain-Limiting energy with Analytic eigensystem
We develop our cubic strain-limiting energy based on the FEM Baraff-Witkin (FBW) constitutive model (Kim, 2020). Inspired by Li et al. (2021) which augments soft membrane energy with a barrier-based strain-limiting term to avoid membrane locking, we aim to develop a barrier-free strain-limiting energy with analytic eigensystems so that expensive backtracking-based line search filtering for the strain limit and numerical eigendecomposition are both avoided. We will first introduce FBW and our model (subsection 5.1), and then derive the analytic eigensystem of our Hessian matrix (subsection 5.2).
5.1. Formulation
In the FBW method (Kim, 2020), the membrane energy is defined as the sum of a stretching term and a shearing term:
(8) $\Psi_{memb}=\Psi_{stretch}+\Psi_{shear}.$
The stretching term $\Psi_{stretch}$ is defined using two orthonormal basis, $\mathbf{n}_{u}=[1,0]^{T}$ and $\mathbf{n}_{v}=[0,1]^{T}$, and based on $I_{5}(\mathbf{F},\mathbf{n})=\mathbf{n}^{T}\mathbf{F}^{T}\mathbf{F}\mathbf{n}$ (Kim et al., 2019), where $\mathbf{F}=[\mathbf{F}_{0}|\mathbf{F}_{1}]\in\rm I\!R^{3x2}$ is the deformation gradient:
(9) $\Psi_{stretch}=\lambda a_{t}(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}-1)^{2}+\lambda a_{t}(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{v})}-1)^{2}.$
Here, $a_{t}$ is the volume weight, i.e. the product of triangle area and the thickness, and $\lambda$ is the stretching stiffness. In this paper, we simply use the same $\lambda$ for both $\mathbf{n}_{u}$ and $\mathbf{n}_{v}$ directions. But note that different stiffnesses can be applied to model more anisotropic behaviors, which will not affect the computation and eigenanalysis.
We follow Kim (2020) and use $\sqrt{I_{5}(\mathbf{F},\mathbf{n})}$ to define our strain-limiting energy, as it effectively represents the stretch factor in the direction $\mathbf{n}$. Directly incorporating $I_{5}$ into the barrier-based strain-limiting model of Li et al. (2021) enables reusing the eigenanalysis from Kim (2020). However, backtracking-based line search filtering, which is not parallel-friendly, is still required since no analytic bound for feasible step sizes is available for the strain. Additionally, because the barrier function diverges rapidly near the strain limit, solver performance becomes highly sensitive to the barrier stiffness, complicating parameter setting. To address this, we apply a cubic penalty function to handle the inequality constraint imposed by the strain limit. This approach allows for slight constraint violations while providing stretching resistance on the same scale as practical settings.
To simplify the explanation, let’s consider the $\mathbf{n}_{u}$ direction in Equation 9 as an example. We construct a $C^{2}$-continuous clamped cubic strain-limiting energy using $I_{5}$:
(10) $\Psi_{SL}^{u}=\lambda^{\prime}a_{t}S(I_{5}(\mathbf{F},\mathbf{n}_{u}))(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}-1)^{3},$
where $\lambda^{\prime}$ is the stiffness and $S$ is a piecewise function defined as
(11) $S(s)=\begin{cases}1,&s>1,\\ 0,&0\leq s\leq 1,\end{cases}$
and so the strain-limiting force is only applied when the cloth is stretched. By similarly defining the cubic term in the $\mathbf{n}_{v}$ direction, we obtain our complete cubic strain-limiting energy
$\Psi_{SL}=\Psi_{SL}^{u}+\Psi_{SL}^{v}.$
Figure 8. Our composed strain energy in the $\mathbf{n}_{u}$ direction. Combined with the membrane energy from Kim (2020), our cubic strain-limiting energy enables using soft compression resistance and stiff stretching resistance to avoid membrane locking and over-elongation issues. Here, $\lambda a_{t}=1$ and $\lambda^{{}^{\prime}}=100\lambda$ are used for visualization.
This energy is well-defined even if the strain limit is exceeded, eliminating the need for line search filtering. Given that the Young’s modulus of real fabrics ranges from $1MPa$ to $10MPa$ (Penava et al., 2014) when the cloth is stretched, setting our $\lambda^{\prime}=5MPa$ provides stretching resistance at the same scale when the cloth approaches the strain limit, e.g., $5\%$. Thus, our model prevents over-elongation in practical cloth manipulation scenarios. For $\Psi_{stretch}$, we can now safely use a small stiffness ($\lambda=0.05MPa$ in all our examples) to avoid membrane locking. Refer to Figure 8 for a plot of our composed membrane energy in the $\mathbf{n}_{u}$ direction. Since the stiffness for the shearing term $\Psi_{shear}$ primarily depends on fabric type and, based on our observations, has minimal impact on the over-elongation or membrane locking issues, we simply set it to 30% of the stretching stiffness and apply the original eigenanalysis from Kim (2020) to compute its PSD Hessian.
5.2. Eigenanalysis
Following Kim et al. (2019); Kim (2020), we can analytically derive the eigensystem of our $I_{5}$-based strain-limiting energy’s Hessian matrix with respect to the deformation gradient $\mathbf{F}$. For simplicity, we will omit $\lambda^{\prime}$ and $a_{t}$ in the following discussion.
The eigenvalues of $I_{5}$-based energies are:
$e_{1}(I_{5})$ $=4I_{5}\frac{\partial^{2}\Psi}{\partial{I_{5}}^{2}}+2\frac{\partial\Psi}{\partial{I_{5}}},$
(12) $e_{2,3}(I_{5})$ $=2\frac{\partial\Psi}{\partial{I_{5}}}.$
Since $\mathbf{n}_{u}$ and $\mathbf{n}_{v}$ are orthogonal, we can substitute $I_{5}(\mathbf{F},\mathbf{n}_{u})$ and $I_{5}(\mathbf{F},\mathbf{n}_{v})$ into subsection 5.2 separately to obtain all six eigenvalues for our $\Psi_{SL}$ when the cloth is stretched ($I_{5}>1$):
(13) $e_{1}(I_{5}(\mathbf{F},\mathbf{n}_{u}))=6(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}-1),$
(14) $e_{2,3}(I_{5}(\mathbf{F},\mathbf{n}_{u}))$ $=3\left(\frac{1}{\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}}+\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{u})}-2\right),$
(15) $e_{4}(I_{5}(\mathbf{F},\mathbf{n}_{v}))=6(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{v})}-1),$
(16) $e_{5,6}(I_{5}(\mathbf{F},\mathbf{n}_{v}))$ $=3\left(\frac{1}{\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{v})}}+\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{v})}-2\right).$
From these equations, we see that all eigenvalues are positive when $I_{5}>1$, indicating that our cubic strain-limiting energy $\Psi_{SL}$ is convex with respect to $\mathbf{F}$ (recall that $\Psi_{SL}=0$ when $I_{5}\leq 1$). Since $\mathbf{F}$ is a linear function of the degrees of freedom $x$, $\Psi_{SL}$ is also convex with respect to $x$.
Thus, we can directly compute the PSD Hessian matrix of $\Psi_{SL}$ in the $\mathbf{n}_{u}$ and $\mathbf{n}_{v}$ directions separately, applying the chain rule with $I_{5}$ as the intermediate variable:
$\mathbf{H}_{SL,*}$ $=3S(I_{5}(\mathbf{F},\mathbf{n}_{*}))\left(1-\frac{1}{\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{*})}}\right)(\sqrt{I_{5}(\mathbf{F},\mathbf{n}_{*})}-1)\mathbf{H}_{*}$
(17) $\,\,+\frac{3S(I_{5}(\mathbf{F},\mathbf{n}_{*}))(I_{5}(\mathbf{F},\mathbf{n}_{*})-1)}{I_{5}(\mathbf{F},\mathbf{n}_{*})^{3/2}}(\mathbf{f}_{*}\mathbf{f}_{*}^{T}).$
Here, $\mathbf{f}_{*}=\text{vec}(\mathbf{F}\mathbf{L}_{*})$, $\mathbf{L}_{*}=\mathbf{n}_{*}(\mathbf{n}_{*})^{T}$, $\mathbf{H}_{*}=\mathbf{L}_{*}\otimes\mathbf{1}_{3\times 3}$, $\otimes$ denotes the tensor product, $\mathbf{I}_{3\times 3}$ is the $3\times 3$ identity matrix, and $\text{vec}(\cdot)$ is the vectorization operator (Kim and Eberle, 2022), with * representing either $u$ or $v$. All of these computations can be efficiently parallelized on the GPU, unlike numerical eigendecomposition.
6. Fast global hessian operations on the GPU
In this section, we introduce our multi-layer reduction strategy for fast Hessian matrix assembly in affine-deformable coupling simulations (subsection 6.1). Using a custom hash-based parallel reduction algorithm (subsection 6.2), our strategy significantly reduces the number of numerical operations and write conflicts, and it enables the development of a memory-efficient symmetric sparse blockwise matrix-vector multiplication method (subsection 6.3) to further boost PCG performance.
6.1. Multi-Layer Reduction for ABD Hessian Assembly
Figure 9. Example of accumulating contact information for reduced affine bodies.
As previously mentioned, in affine body simulation, the main bottleneck lies in transforming and accumulating the contact Hessian matrices (collision and friction) from mesh node DOFs to the affine body DOFs within the global Hessian matrix. For contact pairs involving nodes belonging to affine bodies, consider the example of contact between two affine bodies. In this case, each $3\times 3$ block $\mathbf{C}(r,c)$ in the contact Hessian, where $r$ and $c$ are node indices, must be mapped to a $12\times 12$ matrix for the affine body pair $\{AF(r),AF(c)\}$ via multiplication with the Jacobian matrices $\{\mathbf{J}_{r},\mathbf{J}_{c}\}$. All resulting $12\times 12$ Hessian matrices for each affine body pair are then accumulated (Lan et al., 2022a; Chen et al., 2022).
However, the number of mapped Hessians typically far exceeds the number of affine body pairs (see Figure 9), leading to significant accumulation overhead, let alone the frequent write conflicts. To address this issue, we first accumulate all $3\times 3$ contact Hessian blocks with the same node index pair $\{r,c\}$ through reduction, then apply a single product with $\{\mathbf{J}_{r},\mathbf{J}_{c}\}$ to transform it into a $12\times 12$ matrix. This process yields a sequence of $12\times 12$ matrices for the blocks of the global Hessian corresponding to the affine body pair $\{AF(r),AF(c)\}$. A second layer of reduction is then used to accumulate these $12\times 12$ Hessians in parallel, significantly reducing write conflicts and the number of values to be accumulated, resulting in faster performance.
6.2. Hash-based Parallel Reduction
To further enhance accumulation efficiency, we developed a parallel reduction technique using a hashing method based on either the node index pair $\{r,c\}$ or the affine body index pair $\{AF(r),AF(c)\}$. Each matrix block is encoded with a 64-bit hash value, where the first index forms the higher 32 bits and the second forms the lower 32 bits. Sorting the hash values enables consecutive memory placement of matrix blocks sharing the same hash (Huang et al., 2019, 2020), facilitating efficient parallel accumulation. The key insight here is to build a custom reduction scheme that is based on registers rather than shared memory and avoid global synchronization by taking advantage of warp-level operations (Zhao et al., 2020; Gao et al., 2018; Huang et al., 2024).
Our algorithm is constructed using basic parallel primitives:
  1. (1)
    Device run-length encoding
  2. (2)
    Device exclusive sum
  3. (3)
    Warp segmented reduction
The first two primitives are standard in parallel computing. For warp-level segmented reduction, we use the CUB intrinsic function cub::WarpReduce(values, tags). Consider a simple example with a warp size of 8: if $\emph{{values}}=[1,1,1,1,1,1,1,1]$ and $\emph{{tags}}=[1,0,0,1,0,0,1,0]$, the output will be $\emph{{outPut}}=[3,\bullet,\bullet,3,\bullet,\bullet,2,\bullet]$, where $\bullet$ indicates invalid entries. Here, the tags array specifies start and end positions for reduction in values, with 1 marking the start. Reduction results (valid entries in outPut) are then written to the result array. The overall procedure, FastHashReduction, is detailed in Algorithm 2. First, the map array is constructed in parallel, mapping duplicate keys to their unique key indices to ensure reduction results are written to the correct locations (lines 1–8). The FastSegmentReduction algorithm (Algorithm 3) then efficiently computes the matrix.
Algorithm 2 FastHashReduction
1:Input:
2:     $K$triangleright an array of sorted hash keys
3:     $V$triangleright an array of values, sorted by the hash keys
4:Output:
5:     $R$triangleright an array of reduced values
6:Method:
7:$UK,NK\leftarrow$ RunLengthEncode($K$) $\triangleright$ get unique keys ($UK$) and their counts ($NK$)
8:$Num\leftarrow$ len(UK)
9:$ESum\leftarrow$ ExclusiveSum($NK$) $\triangleright$ get prefix sum, starting from 0
10:// assign the unique key index to $O$ for each key in $K$:
11:$P\leftarrow$ Zeros(len(K))
12:for $tid=0,1,...,Num-1$ do in parallel
13:     $P[ESum[tid]+NK[tid]-1]\leftarrow 1$
14:end for
15:$O\leftarrow$ ExclusiveSum($P$) $\triangleright$ map each $K$ to unique key
16:$R\leftarrow$ FastSegmentReduction($O$,$V$)
FastHashReduction
Our FastHashReduction algorithm is designed to handle various value types. For ABD Hessian matrices, each value is a $12\times 12$ matrix, while for FEM Hessian matrices, each value is a $3\times 3$ matrix. In a hybrid FEM-ABD model, there are four types of Hessian matrix blocks: ABD-ABD ($12\times 12$), FEM-FEM ($3\times 3$), ABD-FEM ($12\times 3$), and FEM-ABD ($3\times 12$). To unify these four block matrix types, we decompose them into multiple $3\times 3$ sub-blocks, resulting in:
  1. (1)
    16 hash keys and 16 sub-blocks of $3\times 3$ matrices for ABD-ABD Hessian matrix blocks.
  2. (2)
    4 hash keys and 4 sub-blocks of $3\times 3$ matrices for ABD-FEM and FEM-ABD Hessian matrix blocks.
The FEM-FEM Hessian matrix blocks remain unchanged. This approach allows us to launch a single, unified CUDA kernel to reduce all types of local Hessians, maximizing parallelism and minimizing branching.
Figure 10. Illustration of FastSegmentReduction. The warp-level reduction is performed within segments, using Array $O$ to identify segment boundaries. For segments spanning multiple warps (B and D), atomic operations are applied to avoid data conflicts, while direct writing is used otherwise.
FastSegmentReduction
In Algorithm 2, we create a mapping array $O$ to associate each key-value pair with its target memory location, as illustrated in Figure 10. Array $O$ is also used to generate tag information for the CUB intrinsic function cub::WarpReduce(values, tags). The tags are determined by comparing adjacent entries in $O$, as described in Algorithm 3. Here, $b^{-}$, $b$, and $b^{+}$ represent the previous, current, and next unique key indices, respectively. The IsValid flag checks whether the GlobalThreadId falls within valid data range. If $b^{-}\neq b$, it marks the start of a new segment. The IsCrossWarp flag indicates if the segment spans multiple warps, necessitating atomic operations to avoid data corruption; otherwise, direct writing is safe. These flags are set per thread, but only the leading thread (head thread) can write to the output array. Therefore, we merge flags across threads in each segment: a head thread is valid if any thread in the segment is valid and marked as cross-warp if any spans multiple warps. This is achieved via segmented warp reduction, using an OR operation for merging (see lines 5-28).
This hash key reduction method efficiently constructs the global Hessian matrix, critical for the subsequent linear system solve and preconditioner computation. Additionally, our Hessian assembly method ensures that $3\times 3$ sub-blocks in each row are sorted in the memory by column index, enabling an efficient SpMV algorithm with warp reduction to enhance PCG solver performance.
Algorithm 3 FastSegmentReduction
1:Input:
2:     $O$triangleright The result in line 8 of 2
3:     $V$triangleright An array of values, sorted by the hash keys
4:Output:
5:     $R$triangleright An array of reduced values, one value for each hash key
6:Method:
7:define BD = BLOCK_DIM
8:define WD = WARP_DIM
9:BlockCount $\leftarrow$ (len(V) + BD - 1) / BD
10:for GlobalThreadId $=$ 00, $1$, …, BlockCount×$\times$×BD-1  do in parallel
11:     // get tags for cub::WarpReduce(values, tags):
12:     LaneId $\leftarrow$ GlobalThreadId % WD
13:     $b^{-},b,b^{+}\leftarrow-1$triangleright $-1$ for invalid value
14:     IsValid, IsHead, IsCrossWarp $\leftarrow$ 0
15:     Value $\leftarrow$ $\mathbf{0}$
16:     if $0<$ GlobalThreadId $<$ len(V) then
17:         $b^{-}\leftarrow O[$GlobalThreadId 1]-1]- 1 ]
18:     end if
19:     if GlobalThreadId $<$ len(V) - 1 then
20:         $b^{+}\leftarrow O[$GlobalThreadId +1]+1]+ 1 ]
21:     end if
22:     if GlobalThreadId $<$ len(V) then
23:         $b\leftarrow O[$GlobalThreadId]]]]
24:         Value $\leftarrow V[$GlobalThreadId]]]]
25:         IsValid $\leftarrow 1$
26:     end if
27:     if LaneId ==0==0= = 0 then
28:         IsHead $\leftarrow 1$, IsCrossWarp $\leftarrow b^{-}==b$
29:     else
30:         IsHead $\leftarrow b^{-}\neq b$
31:         if LaneId $=$ WARP_DIM $-1$ then
32:              IsCrossWarp $\leftarrow b^{+}==b$
33:         end if
34:     end if
35:     Tuple $\leftarrow$ (IsValid, IsCrossWarp)
36:     Tuple $\leftarrow$ cub::WarpReduce(Tuple, IsHead)
37:     Value $\leftarrow$ cub::WarpReduce(Value, IsHead)
38:     // write the result to the corresponding memory location:
39:     if IsHead and IsValid then
40:         if IsCrossWarp then
41:              AtomicAdd($R[b]$, Value)
42:         else
43:              $R[b]\leftarrow$ Value
44:         end if
45:     end if
46:end for
6.3. Symmetric Reduce-By-Key (SRBK) SpMV
Sparse matrix-vector multiply (SpMV) is a frequently called subroutine in scientific computing, where a conjugate gradient solver, for example, may invoke hundreds or thousands of SpMV operations for each linear solve. Thus, speeding up SpMV is critical for improving the overall performance. The NVIDIA cuSparse library’s SpMV on Block Sparse Row (BSR) format is the fastest for general sparse matrices, but it lacks support for symmetric matrices. Since the SpMV algorithm is memory-bound, we aim to leverage symmetry to reduce memory access and achieve significant speedup.
We propose a symmetric Reduce-By-Key (SRBK) SpMV method (Algorithm 4) that is approximately 2×2$\times$2 × faster than cuSparse BSR. Our algorithm works on a format called Sorted Symmetric Block Coordinates, storing only the diagonal and upper triangular blocks of the matrix. All block row and column indices are sorted, a free lunch from our FastHashReduction algorithm for assembling the matrix.
Algorithm 4 Symmetric Reduce-By-Key (SRBK) SpMV
1:Input:
2:     $\mathbf{Rid}$, $\mathbf{Cid}$, $\mathbf{AU}$triangleright block row index, block column index, and value of upper-triangular and diagonal blocks
3:     $\mathbf{V}_{input}$triangleright SpMV input vector
4:Output:
5:     $\mathbf{V}_{output}$triangleright SpMV output vector
6:Method:
7:define BD = BLOCK_DIM
8:define WD = WARP_DIM
9:BlockCount $\leftarrow$ (len($\mathbf{AU}$) + BD - 1) / BD
10:for GlobalThreadId $=$ 00, $1$, …, BlockCount×$\times$×BD11-1- 1  do in parallel
11:     Run line 5-29 in 3, replacing $O$ with $\mathbf{Rid}$
12:     if IsValid then $\triangleright$ GlobalThreadId is within valid data range
13:         j $\leftarrow\mathbf{Cid}[$GlobalThreadId]]]]
14:         $\mathbf{H}\leftarrow\mathbf{AU}[$GlobalThreadId]]]]
15:         Value $\leftarrow\mathbf{H}\cdot\mathbf{V}_{input}[j]$triangleright multiply with upper-triangular and diagonal blocks
16:         if $b\neq j$ then $\triangleright$ If not diagonal
17:              AtomicAdd($\mathbf{V}_{output}[j]$, $\mathbf{H}^{T}\cdot\mathbf{V}_{input}[b]$) $\triangleright$ multiply with lower-triangular blocks
18:         end if
19:     end if
20:     Run line 28-37 in 3, replacing $R$ with $\mathbf{V}_{output}$
21:end for
Our SRBK SpMV and FastSegmentReduction algorithms have a similar structure, both using a ”Reduction By Key” parallel primitive. Sorted row indices serve as hash keys, while the product of the block matrix and the input vector represents the value to be reduced (lines 5-9). The main difference is in how we access the lower triangular part of the matrix, where we atomically add the multiplication results to the output vector (line 11). The data from the lower triangular part is directly transposed from the upper part, reducing global memory access by nearly half, and thus significantly enhancing performance.
7. Experiments and Analysis
We present our evaluation and results in this section. The experiments were conducted with a 24-core Intel Core i9 13900KF CPU with 64GB of RAM, and an NVIDIA RTX 4090 GPU with 24GB of RAM. For the simulations, we use modified PCG, with a relative error tolerance of $10^{-4}$, which has shown to be sufficient for IPC (Huang et al., 2024).
Our experiments include ablation and comparative studies on preconditioners, strain-limiting energies, SpMV, and GPU IPC methods (subsection 7.1), followed by stress tests (subsection 7.2). The simulation statistics are presented in Table 1, Table 2, and Table 3. Regarding alternative GPU IPC methods, we compared our system against GIPC (Huang et al., 2024), using the same linear solver and termination criteria as ours, but with hybrid block-diagonal or GPU MAS (Wu et al., 2022) preconditioner. Additionally, we compared our system to the GPU-based ABD method (Lan et al., 2022a), where only CHOLMOD compiled with Intel MKL LAPACK and BLAS is running on the CPU. Finally, we conducted a comparison with the state-of-the-art affine-deformable coupling IPC framework, ZeMa (Du et al., 2024), which also uses the CPU-based CHOLMOD solver, while leveraging GPU acceleration for other components. All simulations were performed using double-precision floating-point arithmetic.
Figure 11. Octopus stack. The simulation uses $\hat{d}=10^{-3}l$, $\rho=1000\,\text{kg}/\text{m}^{3}$, $\Delta t=5\times 10^{-3}\,\text{s}$, and a Newton tolerance of $10^{-2}l\Delta t$. This test case validates the performance improvement of our connectivity-enhanced MAS preconditioner, which achieves a 1.8×1.8$\times$1.8 × speedup in linear solve.
7.1. Ablation and Comparative Studies
Preconditioner.
To evaluate the performance of our connectivity-enhanced MAS preconditioner, we construct several scenes with different geometric structures (see Figure 11, Figure 13, Figure 19, and Figure 20). The timing breakdown and CG iteration counts are reported in Table 1. For a fair comparison, we implemented alternative preconditioners into our simulation framework, ensuring that all other components remain the same. The block-diagonal preconditioner was also tested as a reference. Our results show that, with more effective aggregation, our method using a subdomain size of 32 provides the most significant improvement in PCG convergence compared to MAS (Wu et al., 2022) (see avg. #cg column in Table 1). Even with smaller subdomain sizes (16), our method converges faster than MAS and performs the best overall, achieving up to 3$\times$ speedup in the PCG solve and up to 2.6$\times$ faster for the full simulation, thanks to more efficient precomputation and preconditioning.
As seen from the similar average CG iterations across resolutions (the rows for Figure 11 in Table 1), our preconditioner maintains the scalability of MAS, resulting in near-linear total timing with respect to resolution. In Figure 13, the ’Ours (16)’ configuration shows only a 1.17$\times$ speedup in convergence, as the volumetric structure of the bunnies results in similar METIS and Morton code reorderings, limiting improvement. However, we still observe a 1.58$\times$ speedup in PCG time and a 1.47$\times$ overall speedup, due to the lower computational cost of our preconditioner compared to MAS.
Table 1. Simulation statistics with different preconditioners. Mesh DOFs: number of vertices (v), tetrahedra (t), and surface triangles (f). PPC, PPR, and CGR represent the time spent on preconditioner pre-computation, preconditioning, and CG iterations, respectively, with PCG = PPC + PPR + CGR. The remaining columns report the time for miscellaneous tasks (misc), total simulation time (timeTot), total CG iterations (#cg), total Newton iterations (#Newton), and the average CG iterations per Newton iteration (avg. #cg). All times are in seconds.
Example Mesh DOFs: v, t, f Preconditioner PPC PPR CGR PCG Misc TimeTot #cg #Newton avg. #cg
Figure 11 $42K,171K,45K$ GPU MAS 38 55 102 195 26 221 6.98e5 2.03e3 344
Block Diagonal 0.06 21 338 359 27 385 2.24e6 2.07e3 1082
Ours (32) 32 25 60 117 26 143 3.45e5 2.02e3 171
Ours (16) 10 29 69 108 26 134 4.41e5 1.99e3 221
Speed Up 3.80$\times$ 1.90$\times$ - 1.81$\times$ - 1.65$\times$ - - 1.51$\times$
Figure 11 $114K,558K,69K$ GPU MAS 105 162 257 524 69 593 9.51e5 2.31e3 412
Block Diagonal 0.16 43 582 625 69 694 2.48e6 2.30e3 1079
Ours (32) 96 79 127 302 69 371 4.37e5 2.31e3 189
Ours (16) 33 60 145 238 69 307 5.20e5 2.31e3 225
Speed Up 3.18$\times$ 2.7$\times$ - 2.20$\times$ - 1.93$\times$ - - 1.84$\times$
Figure 13 $203K,1122K,50K$ GPU MAS 152 188 229 569 87 656 6.23e5 1.76e3 354
Block Diagonal 0.18 63 918 981 88 1069 2.61e6 1.77e3 1474
Ours (32) 134 137 180 451 87 538 4.54e5 1.76e3 258
Ours (16) 42 108 210 360 87 447 5.30e5 1.76e3 301
Speed Up 3.61$\times$ 1.74$\times$ - 1.58$\times$ - 1.47$\times$ - - 1.17$\times$
Figure 19 $123K,330K,211K$ GPU MAS 371 836 1029 2236 217 2453 4.38e6 7.93e3 552
Block Diagonal 0.44 137 1920 2057 216 2273 8.84e6 7.90e3 1119
Ours (32) 354 196 298 848 215 1063 9.98e5 7.86e3 127
Ours (16) 76 215 439 730 215 945 1.81e6 7.85e3 231
Speed Up 4.88$\times$ 3.89$\times$ - 3.06$\times$ - 2.60$\times$ - - 2.39$\times$
Figure 20 $79K,298K,110K$ GPU MAS 86 193 274 553 42 595 1.42e6 1.62e3 877
Block Diagonal 0.08 57 798 855 43 898 5.09e6 1.64e3 3105
Ours (32) 47 89 124 260 42 302 6.84e5 1.63e3 420
Ours (16) 12 68 136 216 41 257 7.53e5 1.62e3 465
Speed Up 7.17$\times$ 2.83$\times$ - 2.56$\times$ - 2.31$\times$ - - 1.89$\times$
(a) $Y=5MPa$, no SL
(b) $Y=0.5MPa$, no SL
(c) $Y=50KPa$, no SL
(d) $Y=50KPa$, with Li et al. (2021)
(e) $Y=50KPa$, with our SL
Figure 12. Membrane locking and strain limiting. The cloth consists of $100K$ vertices and $200K$ triangles, simulated with FEM, while the bunny is simulated using ABD. The simulation parameters are: $\Delta t=0.01\,\text{s}$, Poisson’s ratio $\nu=0.49$, density $\rho=200\,\text{kg}/\text{m}^{3}$, cloth thickness $10^{-3}\,\text{m}$, friction coefficient $\mu=0.1$, Newton tolerance $\varepsilon_{d}=10^{-2}l\Delta t$, membrane Young’s modulus $Y=\{5\times 10^{6},5\times 10^{5},5\times 10^{4}\}\,\text{Pa}$, and distance threshold $\hat{d}=10^{-3}l$, where $l$ is the scene diagonal length. Li et al. (2021) (d) and our method (e) simulate fine wrinkling details without membrane locking (a, b) or over-elongation (c), with our method being 1.24×1.24$\times$1.24 × faster (Table 2).
Strain limiting.
Figure 12 presents a cloth simulation example using FBW (Kim, 2020) with and without barrier-based (Li et al., 2021) or our cubic strain-limiting energy. The cloth consists of 100K nodes and is tested with various membrane stiffnesses. In cloth simulation, avoiding excessive stretching is crucial, as it can cause unrealistic, overly-elastic behavior. To prevent this, a large Young’s modulus (e.g., $5MPa$), within the range of real cloth parameters (Penava et al., 2014), is needed. However, such stiffness can lead to membrane locking in even medium-resolution meshes, resulting in sharp creases and artificial plastic appearances (12(a)). Reducing the stiffness to $0.5MPa$ mitigates membrane locking but visible artifacts remain (12(b)). Further reducing the stiffness to $50KPa$ smooths wrinkles but causes over-elongation (12(c)). Our cubic strain-limiting energy provides sufficient stretching resistance as in real fabrics, avoiding excessive stretching and capturing smooth, realistic wrinkles with a small membrane stiffness of $50KPa$ for compression (12(e), and supplemental video).
We also compare with the barrier-based strain-limiting method in CIPC using their default settings: a $1.1$ strain limit and $1$ activation threshold. We use a $1KPa$ stiffness for the barrier term, as smaller or larger values both lead to worse conditioned systems. With similar result quality (12(d)), our method is 1.24×1.24$\times$1.24 × faster (Table 2), with lower costs in local Hessian construction and line search (see Table 3 for a more detailed breakdown). While our method performs well for practical cloth manipulation scenarios, it does not guarantee strain limit satisfaction in extreme cases like the barrier-based method. Thus, we further compare their behavior in a stress test (Figure 22), detailed later.
Table 2. Simulation statistics for Figure 12. Columns show total time for linear solve (PCG) and miscellaneous tasks (Misc), total simulation time (TimeTot), total PCG iterations (#cg), total Newton iterations (#Newton), and average PCG iterations per Newton iteration (#cg per iter). All simulations use the same PCG relative tolerance of $10^{-4}$ with our connectivity-enhanced MAS preconditioner.
Young’s Modulus ($Pa$) PCG Misc TimeTot #cg #Newton avg. #cg per iter
$5MPa$, no SL (12(a)) 2.91e2 1.34e2 4.25e2 7.07e5 4.03e3 175
$0.5MPa$, no SL (12(b)) 2.31e2 1.61e2 3.92e2 4.12e5 4.49e3 92
$50KPa$, no SL (12(c)) 1.26e2 1.34e2 2.60e2 1.66e5 3.63e3 46
$50KPa$ + CIPC SL (12(d)) 1.55e2 2.02e2 3.57e2 2.24e5 3.99e3 56
$50KPa$ + our SL (12(e)) 1.45e2 1.43e2 2.88e2 2.06e5 3.76e3 55
Figure 13. Stiff and soft bunnies. The top bunny is nearly rigid (Young’s modulus $10^{8}$ Pa), and the bottom red one is deformable (Young’s modulus $10^{6}$ Pa). Each bunny consists of $101K$ vertices and $561K$ tetrahedra. The simulation uses $\hat{d}=5\times 10^{-4}l$, $\rho=1000\,\text{kg}/\text{m}^{3}$, $\Delta t=5\times 10^{-3}\,\text{s}$, and a Newton tolerance of $10^{-2}l\Delta t$. This test compares GIPC, ZeMa, and our method in hybrid scenarios. Our method shows a 4$\times$ speedup when simulating both bunnies using FEM and a 10$\times$ speedup when using ABD+FEM.
Global Hessian assembly
As illustrated in Figure 14, we use the wrecking ball simulation (Figure 18) to conduct an ablation study on our multi-layer reduction strategy (Strategy C) for global Hessian matrix assembly, comparing it against single-layer reduction (Strategy B) and pure atomic operations (Strategy A). In both Strategies A and B, the local Hessian matrices of all contact pairs are first mapped to the affine body DOFs by multiplying the Jacobian matrices. Strategy B then performs a reduction to accumulate these matrices into the global Hessian, while Strategy A uses atomic operations directly for accumulation. To ensure a fair comparison, we also sort the value array prior to atomic accumulation. This step optimizes the distribution of the value array, enabling aggregation optimization in the CUDA intrinsic function atomicAdd, thereby achieving the best possible performance. Even with this optimization, pure atomic operations remain slightly slower than single-layer reduction, as shown in the comparison of Strategies A and B in Figure 14. Our multi-layer reduction strategy achieves an additional nearly 5×5$\times$5 × speedup over single-layer reduction.
Reduce by Key
We compare our FastHashReduction method with the CUB intrinsic function cub::DeviceReduce::ReduceByKey and plot our speedup relative to the number of values being accumulated (Figure 15). To keep the total usage of GPU memory unchanged, the reduction is performed on 16M $3\times 3$ matrices, and we adjust the number of matrices with the same row or column indices to control the workload. Our method shows superior performance across a wide range of accumulation workload, from $1$ to $2^{16}$, achieving a maximum speedup of approximately 1.55×1.$5$ at $2^{11}$. In FEM simulations, the average number of values with the same row or column indices typically remains below $2^{7}$, where our method achieves a consistent speedup of $1.2$ to 1.4×1.4$\times$1.4 ×. Additionally, we observe that cub::DeviceReduce::ReduceByKey fails to compile when reducing the $12\times 12$ affine body Hessian matrix due to shared memory limitations. In contrast, our FastHashReduction method remains effective by reusing warp registers, leveraging the element-independent nature of matrix summation.
SpMV
We compare the performance of different SpMV methods using the stiff and soft bunnies example shown in Figure 13, simulating both bunnies with FEM. As shown in Figure 16, the BSR SpMV (1$\times$) from cuSparse [^1] https://developer.nvidia.com/cusparse is treated as a baseline. Without matrix sorting, the Triplet SpMV (0.16$\times$) implemented by atomically adding the matrix-vector products and the MatrixFree SpMV (0.18$\times$) from GIPC both suffer from an atomic operation flood, which undermines the parallelism. On the other hand, sorting-based methods have a significant speedup, benefiting from the reduction of atomic operations. The traditional CSR SpMV (0.75$\times$) from cuSparse has an almost 5$\times$ speedup compared to the unsorted methods. The Block Coordinates (BCOO) SpMV from MUDA [^2] https://github.com/MuGdxy/muda and the BSR SpMV from cuSparse further improves the performance by applying blockwise matrix-vector multiplication. Our SRBK SpMV method (1.85×1.85$\times$1.85 ×) further mitigates the memory-bound issue of BSR SpMV and boosts the performance by almost 2×2$\times$2 ×.
To evaluate the scalability of our method, we compare its speedup factor to BSR on matrices with varying numbers of non-zero $3\times 3$ blocks per row. As shown in Figure 17, the speedup factor increases from 0.93×0.93$\times$0.93 × to 1.9×1.9$\times$1.9 × as the number of non-zero blocks approaches $2^{3}$ from 0. For average non-zero block counts between $2^{3}$ and $2^{6}$ – the typical range for FEM global systems – the speedup factor fluctuates between 1.8×1.8$\times$1.8 × and 2×2$\times$2 ×, which demonstrates strong scalability.
Figure 14. Ablation study on multi-layer reduction for Hessian assembly. We compare our multi-layer reduction strategy (C) against single-layer reduction (B) and pure atomic operations (A) in the wrecking ball simulation (Figure 18). Despite sorting the value array for optimal atomicAdd performance, using atomic operations is slightly slower than single-layer reduction, and ours achieves nearly 5×5$\times$5 × speedup over Strategy B.
Figure 15. Speedup of FastHashReduction vs. CUB ReduceByKey. Speedup factor is shown relative to the average number of matrices sharing the same row or column indices. A peak speedup of 1.55×1.$5$ occurs around $2^{11}$. For typical duplication levels in FEM simulations (up to $2^{7}$), FastHashReduction maintains a consistent speedup between 1.2×1.2$\times$1.2 × and 1.4×1.4$\times$1.4 ×.
Figure 16. Timing comparison of SpMV methods. The BSR method from cuSparse serves as the baseline. Our method (SRBK) achieves a maximum speedup of 1.85$\times$ over the baseline.
Figure 17. Scalability of our SRBK SpMV. Using the BSR from cuSparse as the baseline, we compare our SpMV’s speedup over BSR on matrices with different numbers of non-zero blocks per row. The speedup increases from 0.93×0.93$\times$0.93 × to 1.9×1.9$\times$1.9 × as non-zero blocks per row approach $2^{3}$. For typical FEM ranges ($2^{3}$$2^{6}$ blocks), speedup stabilizes between 1.8×1.8$\times$1.8 × and 2×2$\times$2 ×, demonstrating our strong scalability.
Figure 18. Wrecking Ball. This example compares GPU ABD (Lan et al., 2022a) and our method using only stiff affine bodies. The simulation parameters are $\hat{d}=10^{-3}l$, $\rho=1000$ kg/m3, $\Delta t=0.01$ s, $\mu=0.05$, static friction tolerance $10^{-3}l$, and Newton tolerance $10^{-2}l\Delta t$. Our method achieves a 2.6$\times$ speedup.
Comparison with GPU ABD
To evaluate the efficiency of our framework in simulating stiff affine bodies, we replicate the wrecking ball example (Figure 18) from ABD (Lan et al., 2022a). In our setup, the sphere and the boxes are assigned the same density, ensuring that the number of contacts remains large as less boxes will be punched out. This configuration makes the problem more challenging for the simulation framework. Our method utilizes PCG to solve the linear systems, using the $12\times 12$ diagonal blocks as the preconditioner. In this scenario, ABD requires 37.2 ms per Newton iteration, whereas our method only takes 14.4 ms, resulting in a 2.6$\times$ speedup.
Table 3. Simulation statistics for all examples. Columns show time for energy gradient and Hessian computation (Hess), linear system solves (LSolve), CCD (CCD), backtracking line search (LineS), miscellaneous tasks (Misc), and the entire simulation (TimeTot), followed by iteration counts, and finally, speedup relative to GIPC. All times are in seconds. Here, we compare our method with GIPC (Huang et al., 2024), ZeMa (Du et al., 2024), and the barrier-based SL (Li et al., 2021). We also show how each of our innovation contributes to the overall speedup.
Mesh DOFs: v, t, f Framework Hess LSolver CCD LineS Misc TimeTot #cg #Newton avg. #cg Speed Up
Figure 11 114K, 558K, 69K GIPC 63 1134 5 13 6 1221 9.41e5 2.30e3 409 1.00$\times$
114K, 558K, 69K Ours (SRBK) 45 524 5 13 6 593 9.51e5 2.31e3 412 2.06$\times$
114K, 558K, 69K Ours (CEMAS+SRBK) 45 238 5 13 6 307 5.20e5 2.31e3 225 3.98$\times$
Figure 12 109K, 80K, 200K GIPC 106 606 20 54 6 792 2.34e5 3.89e3 60 1.00$\times$
109K, 80K, 200K Ours (SRBK) 75 489 20 55 6 645 2.37e5 3.86e3 61 1.23$\times$
109K, 80K, 200K Ours (CEMAS+SRBK) 75 184 20 55 6 340 2.61e5 3.95e3 66 2.33$\times$
109K, 80K, 200K CIPC (CEMAS+SRBK) 104 202 21 91 6 424 2.89e5 4.20e3 69 -
100K, 0, 200K Ours (ABD+CEMAS+SRBK) 67 145 19 51 6 288 2.06e5 3.76e3 55 2.75$\times$
Figure 13 203K, 1122K, 50K GIPC 90 1481 4 14 7 1596 6.23e5 1.76e3 354 1.00$\times$
114K, 561K, 50K ZeMa 42 54810 2.7 9 7 54871 - 1.74e3 - 0.03$\times$
203K, 1122K, 50K Ours (SRBK) 63 569 3 14 7 656 6.23e5 1.76e3 354 2.43$\times$
203K, 1122K, 50K Ours (CEMAS+SRBK) 63 360 3 14 7 447 5.30e5 1.76e3 301 3.57$\times$
114K, 561K, 50K Ours (ABD+CEMAS+SRBK) 37 104 2.7 8.5 7 159 1.73e5 1.73e3 100 10.03$\times$
Figure 19 123K, 330K, 211K GIPC 137 3910 35 83 8 4173 4.31e6 7.87e3 548 1.00$\times$
123K, 330K, 211K Ours (SRBK) 91 2236 35 83 8 2453 4.38e6 7.93e3 552 1.70$\times$
123K, 330K, 211K Ours (CEMAS+SRBK) 90 730 35 82 8 945 1.81e6 7.85e3 231 4.42$\times$
Figure 20 79K, 298K, 110K GIPC 24 1080 5 13 8 1130 1.42e6 1.62e3 877 1.00×1.$0$
79K, 298K, 110K Ours (SRBK) 16 553 5 13 8 595 1.42e6 1.62e3 877 1.90$\times$
79K, 298K, 110K Ours (CEMAS+SRBK) 16 216 5 12 8 257 7.53e5 1.62e3 465 4.40$\times$
78K, 290K, 110K Ours (ABD+CEMAS+SRBK) 14 148 4 8 7 181 4.50e5 1.41e3 319 6.24$\times$
Figure 21 70K, 241K, 106K GIPC 692 10686 114 637 14 12143 7.30e6 1.58e4 462 1.00$\times$
70K, 241K, 106K Ours (SRBK) 460 4511 113 594 14 5692 7.58e6 1.58e4 480 2.13$\times$
70K, 241K, 106K Ours (CEMAS+SRBK) 452 1547 112 597 14 2722 3.37e6 1.57e4 214 4.46$\times$
53K, 133K, 106K Ours (ABD+CEMAS+SRBK) 407 696 103 535 12 1753 1.48e6 1.46e4 101 6.93$\times$
Figure 22 105K, 10K, 202K GIPC 210 2473 54 211 9 2957 3.10e6 9.43e3 329 1.00$\times$
105K, 10K, 202K Ours (SRBK) 142 762 51 205 9 1069 3.13e6 9.27e3 338 2.76$\times$
105K, 10K, 202K Ours (CEMAS+SRBK) 145 451 52 207 9 864 8.21e5 9.50e3 86 3.42$\times$
105K, 10K, 202K CIPC (CEMAS+SRBK) 219 812 59 329 10 1429 1.75e6 9.69e3 180 -
105K, 0, 202K Ours (ABD+CEMAS+SRBK) 161 468 47 178 9 863 7.78e5 8.51e3 91 3.42$\times$
Comparison with GIPC and ZeMa on hybrid scenarios.
Here, we compare our method with GIPC (Huang et al., 2024) and ZeMa (Du et al., 2024) on the stiff and soft bunnies example (Figure 13). In this simulation, the Young’s modulus is set to $10^{6}$ Pa for the bottom bunny and $10^{8}$ Pa for the top bunny. Our framework can treat the top bunny as a stiff affine body and the bottom bunny with FEM, while GIPC needs to simulate both using FEM. Results in Table 3 indicate that our coupling framework achieves a 10$\times$ overall speedup compared to GIPC. To analyze the sources of this acceleration, we applied our contributions incrementally. ’Ours (SRBK)’ uses only our SRBK method, matching other components to GIPC, which already yields a 2.43×2.43$\times$2.43 × speedup with more efficient Hessian assembly and SpMV. ’Ours (CEMAS+SRBK)’ further adds connectivity-enhanced (CE) MAS preconditioner to ’Ours (SRBK)’, retaining other components, and shows an additional 1.47×1.47$\times$1.47 × overall speedup due to faster PCG convergence enabled by CEMAS, resulting in a 3.57×3.57$\times$3.57 × improvement compared to GIPC. Note that in both ’Ours (SRBK)’ and ’Ours (CEMAS+SRBK)’ cases, we simulate both bunnies using FEM. ZeMa performs the worst in this case, as direct solvers are not efficient for simulations with a large number of DOFs. Thus, for general IPC simulations, GIPC remains the best alternative. Therefore, in subsequent experiments, we compare our methods only against GIPC.
Figure 19. Furry ball.This example compares GIPC (Huang et al., 2024) and our method. The simulation parameters are $\hat{d}=4\times 10^{-4}l$, $\rho=1000$ kg/m3, $Y=1MPa$, $\nu=0.49$, $\Delta t=0.01$ s, and $\varepsilon_{d}=10^{-2}l\Delta t$. Our method achieves a 4.42$\times$ speedup.
Figure 20. Board House. This example compares GIPC (Huang et al., 2024) with our method. Simulation parameters are $\hat{d}=1\times 10^{-4}l$, $\rho=1000\,\text{kg/m}^{3}$, $Y=0.1\,\text{GPa}$, $\nu=0.49$, $\Delta t=0.01\,\text{s}$, and $\varepsilon_{d}=10^{-2}l\Delta t$. Our method achieves a 4.4$\times$ speedup using FEM and 6.24$\times$ using ABD.
We further compare our method with GIPC on the examples shown in Figure 11, Figure 12, Figure 19, Figure 20, Figure 22, and Figure 21, with statistics reported in Table 3. Generally, ’Ours (SRBK)’ achieves a 1.23×1.23$\times$1.23 ×2.76×2.76$\times$2.76 × speedup over GIPC. Adding our CEMAS preconditioner further improves performance, providing an additional 1.24×1.24$\times$1.24 ×2.6×2.6$\times$2.6 × speedup as shown by ’Ours (CEMAS+SRBK)’, leading to an overall improvement of 2.33×2.$3$4.46×4.46$\times$4.46 × compared to GIPC. Notably, this is achieved without approximating stiff materials as affine bodies, which can further boost performance up to 2.75×2.75$\times$2.75 ×10×10$\times$10 × faster than GIPC. In Figure 12, comparing ’Ours (SRBK)’ and ’Ours (CEMAS+SRBK)’ reveals that while CEMAS does not improve PCG convergence due to similar ordering with Morton code sorting in GIPC, it significantly reduces PCG solver time due to lower construction and preconditioning costs. Here, PCG costs represent a smaller portion of total timing than in other examples due to fewer PCG iterations required, resulting in a modest 1.23×1.23$\times$1.23 × speedup with SRBK alone. We will discuss Figure 21 and Figure 22 in detail in the next subsection on stress tests.
7.2. Stress Tests
Mat-cloth twist
To further validate the robustness and efficiency of our framework, we perform a simulation coupling volumetric solids, cloth, and stiff affine bodies, as shown in Figure 21. The Young’s modulus of the cloth and the mat are $50KPa$ and $0.5MPa$ respectively. They both have a Poisson’s ratio of 0.49 and the friction coefficient of 0.4. We twist the cloth and mat by rotating their left and right sides for 2 rounds in 10 seconds, wrapping a stiff affine ellipsoid between them, creating extreme stretching, bending, and frictional contacts. The maximum number of contact pairs is 567K, averaging 351K per time step. The full simulation completed in 29 minutes for 1000 time steps (see the detailed time breakdown in Table 3). Our framework exhibit a 6.93$\times$ speedup compared to GIPC in this case.
Figure 21. Mat-cloth twist. Stress test on extreme deformation with cloth, volumetric mat, and a stiff affine ellipsoid (inside). Our method simulates this stiff challenging scenario efficiently with 1.8s per time step ($\Delta t=0.01s$), achieving a 6.9×6.9$\times$6.9 × speedup compared to GIPC.
Box Pile
[…]
In this example, we observe significantly faster PCG convergence with a diagonal preconditioner than with the MAS preconditioner in GIPC. Thus, we apply the diagonal preconditioner for both GIPC and ’Ours (SRBK)’. This difference arises because Morton code-based reordering may spread nodes from the same box into different groups. Without connections between boxes, nodes from different boxes in the same group cannot be effectively merged, slowing down MAS performance. In contrast, our CEMAS minimizes inter-group connections, effectively grouping nodes from the same box, enabling faster PCG convergence as shown in ’Ours (CEMAS+SRBK)’. Interestingly, approximating the boxes as stiff affine bodies does not significantly improve performance in this scenario, as each box’s small DOF count allows efficient computation with pure FEM.
Figure 22. Box pile. $8\times 30\times 8$ boxes dropped onto a high-resolution cloth with 100K nodes, showing our capability of efficiently simulating high-speed impacts, rapidly changing contacts, and fine wrinkling details in the same scene at 3.45s per time step.
Figure 23. Diagonal view of the London bus. The yellow gear acts as the motor, driving the bus forward through rolling friction between the wheels and the ground. The wheels are simulated as rubber using FEM, while all other parts are modeled as stiff affine bodies. Detailed simulation setup is provided in Table 4.
3D puzzle bus simulation
We digitalize a 3D puzzle, the London Bus from Wooden City® (Figure 23). We extract the part contours in the illustrations of the puzzle instructions to create the 2D sketches in AutoCAD®. The 3D parts are created and assembled in SolidWorks ®. Our eye-norm modeling accuracy of the 3D parts, as well as the rough assembling accuracy, is so far from the real-world vehicle manufacture standard that the gear teeth are not even well aligned, leading to a poor dynamic balance condition on the high-speed rotating parts. Thus, it is challenging to predict the behavior of the bus using the traditional analytic motion analysis (e.g., Mechanical Linkage Analysis), but our framework can still simulate it robustly with the yellow gear acting as the motor (see our supplemental document for details). In fact, our simulator robustly captures the realistic non-smooth motion caused by potential interference in the low-accuracy gear system and self-adjustment of motor speed (averaging $3m/s$) relative to resistance (see Figure 1 and our supplemental video).
Figure 24. Timing breakdown of the London bus simulation. The outer ring shows the high-level breakdown into Gradient/Hessian (21.2%), Linear solve (48.4%), Line search (5.8%), and Misc (24.6%). The inner ring provides a more detailed breakdown.
A timing breakdown of the simulation is shown in Figure 24. Since the bus is entirely driven by frictional contact forces between the FEM wheels and the ground, resulting in an average of 118K rapidly changing contact pairs, the scene complexity leads to a large number of Newton iterations. The median number of Newton iterations per time step is 29, with a maximum of 786. Despite this challenging scenario, our simulator achieves an average of 1.56 seconds per frame. Detailed simulation setup is provided in Table 4.
Table 4. Simulation setup of the London Bus. FEM Y/P refers to the Young’s modulus and Poisson ratio; Gear fr. and Wheel fr. mean the friction coefficient of gear-gear contact and wheel-ground contact, respectively.
Mesh: v, t, f $\Delta t$ FEM Y/P FEM density ABD stiffness ABD density
$(191K,676K,265K)$ $5\times 10^{-3}s$ $0.497MPa/$ 0.49 $500kg/m^{3}$ $100MPa$ $600kg/m^{3}$
FEM: v,t Gear fr. Wheel fr. Relative $\hat{d}$ Bus Dimensions
$(31K,121K)$ $1\times 10^{-3}$ $0.99$ $5\times 10^{-4}$ $0.7m\times 1.2m\times 2.3m$
8. Conclusion and Discussion
In this paper, we presented a novel, fully GPU-optimized Incremental Potential Contact (IPC) framework capable of simulating materials of a wide range of stiffness with consistent high performance and scalability. Our framework introduces several significant advancements over previous methods, achieving up to 10$\times$ speedup compared to the state-of-the-art GIPC. Our contributions include: 1) A novel connectivity-enhanced MAS preconditioner optimized for the GPU, which significantly improves convergence at a lower preconditioning cost; 2) A cubic strain-limiting energy with an analytic eigensystem, designed to simulate stiff membranes such as cloth without membrane locking; 3) An innovative hash-based multi-layer reduction strategy, enabling fast matrix assembly for efficient affine-deformable coupling. Our framework’s robust performance is demonstrated through extensive benchmark studies and rigorous performance analyses, showing superior results in various scenarios, including soft, stiff, and hybrid simulations. The framework efficiently handles high resolution, large deformations, and high-speed impacts, maintaining accuracy and robustness. We believe our advancements address critical challenges in IPC simulation, providing a scalable and high-performance solution for complex frictional contact behaviors. Our comprehensive approach not only improves the efficiency of IPC but also retains the method’s reliability and accuracy, making it a valuable tool for future research and applications in a variety of fields.
Limitations and future work
Our preconditioner was designed under the assumption that the mesh connectivity would not change during the simulation. We precompute the partition when loading the meshes, which means our current design is not suitable for simulations with adaptive mesh refinement and/or fracture mechanics, where the mesh topology may change in any frame. Additionally, the main performance bottleneck of IPC comes from two main factors: the expensive computational cost of each Newton iteration and the large number of Newton iterations required in cases with challenging contact scenarios. Our preconditioner design and GPU optimizations have primarily focused on minimizing the cost of each Newton iteration. However, the slow convergence of the optimization in challenging cases can still result in slower overall performance compared to simulation methods that trade accuracy for efficiency. Therefore, it is very meaningful to keep investigating better preconditioners that can effectively handle simulations with mesh topology changes and to explore novel nonlinear solvers or discretization strategies that can significantly improve the convergence of the nonlinear solver.